CaF: ALL SPECTRA AND ALL DYNAMICS

\underline{R. W. FIELD}, J. J. KAY, S. L. COY, V. S. PETROVIĆ, S. N. ALTUNATA, and B. M. WONG, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139; Ch. JUN-GEN, Laboratoire Aim Cotton du CNRS, Universit de Paris Sud, F-91405 Orsay, France.

More than 1,000 CaF rovibronic levels, spanning the range n^{∗} = 5, v = 0 (42, 500 cm^{−1}) to n^{∗} = 20, v = 1 (47, 500 cm^{−1}), and sampling ℓ = 0(s) − 5(h) are fitted to a quantum defect model. This model consists of two currently separate pieces: the core-penetrating s, p, d, f series are fitted to the elements of an internuclear distance (R) and energy (E) dependent quantum defect matrix, μ(R, E); the core-nonpenetrating series are described by multipole (μ, Q, O) moments and polarizabilities (α, γ). The values of the fitted parameters agree well with ab initio computed values. Although the fitted parameters permit calculation of all spectra and all dynamics of CaF, the physical meanings of the parameters remain obscure.